### v1

@@ A.3 Proof of Proposition 2
Proof of Theorem 4.1 Since $\mathcal{M}$ satisfies Assumption 1 and $h$ is smooth, we can take a neighborhood $\mathcal{B}_{p}\subseteq\mathbb{R}^{q}$ of $X$ such that $c_{1}$ and $h$ are $L_{c}$- and $L_{h}$-Lipschitz on $\mathcal{B}_{p}$, respectively, and $\operatorname{dist}({Y},\mathcal{M})\leq\rho_{\mathcal{M}}\|c_{1}({Y})\|_{2}$ holds for a constant $\rho_{\mathcal{M}}>0$ and any ${Y}\in\mathcal{B}_{p}$ (see Lemma 4). Therefore, we can find a neighborhood $\mathcal{B}\subseteq\mathcal{B}_{p}$ such that given any $Y\in\mathcal{B}$, there exists $Y_{p}\in\mathcal{M}\cap\mathcal{B}_{p}$ such that $\|Y-Y_{p}\|_{2}=\operatorname{dist}(Y,\mathcal{M})$.
@@
where we employ the the triangle inequality $\|h(Y_{p})\|_{2}-\|h(Y)\|_{2}\leq\|h(Y_{p})-h(Y)\|_{2}$, and the Lipschitz continuity of $\phi$ and $h$. Therefore, setting $\rho=\max\{\rho_{\mathcal{M}}(1+L_{\phi}L_{h}C\rho_{\overline{\mathcal{K}}}),\,L_{\phi}C\rho_{\overline{\mathcal{K}}}\}$ concludes that $\operatorname{dist}(Y,\mathcal{M}\cap\mathcal{K})\leq\rho\|(c_{1}(Y),h(Y))\|_{2}$ for all $Y\in\mathcal{B}$, which together with the locally Lipschitz property of $(c_{1},h)$, verifies that $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1. Applying Theorem 1 to $\mathcal{M}\cap\mathcal{K}=\{\tilde{X}\in\mathbb{R}^{q}\mid c_{1}(\tilde{X})=0,\,h(\tilde{X})=0\}$ gives
@@
$=\{\eta\in\mathbb{R}^{q}\mid c_{1}^{\prime\prime}(X;\eta,\zeta)=0\}\cap\{\eta\in\mathbb{R}^{q}\mid h^{\prime\prime}(X;\eta,\zeta)=0\}$
@@ C.3 $\mathcal{H}$ as the product of hyperbolic manifolds
$\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathbb{H}^{n}_{m-1})=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid\bar{h}(X)=0\}\cap\mathcal{Z},$ (C.15)
@@ D.2.1 $\mathcal{U}$ as a Frobenius sphere
where we use the triangle inequality $|\|\tilde{X}_{P}\|_{\mathrm{F}}-1|\leq|\|\tilde{X}\|_{\mathrm{F}}-1|+|\|\tilde{X}\|_{\mathrm{F}}-\|\tilde{X}_{P}\|_{\mathrm{F}}|\leq|\|\tilde{X}\|_{\mathrm{F}}-1|+\|\tilde{X}-\tilde{X}_{P}\|_{\mathrm{F}}$. Consequently, we can confirm the error bound property of $\mathcal{S}_{j}\cap\mathcal{U}$ by incorporating (D.6) and $|\|\tilde{X}\|_{\mathrm{F}}-1|\leq|h(X)|$ into (D.15). Applying Theorem 3.1 to $\mathcal{S}_{j}\cap\mathcal{U}$ and taking into account the rule (2.4) show that
@@ D.3.1 $\mathcal{U}$ as an affine set
Appendix E Proof of Lemma 3
@@
$U_{i}\ \tilde{U}_{i\bot}\ \breve{U}_{i\bot}\ \tilde{U}_{iY\bot}\ U_{iY}]\in\mathcal{O}(m),\qquad[V_{i}\ \tilde{V}_{i\bot}\ \breve{V}_{i\bot}\ \tilde{V}_{iY\bot}\ V_{iY}]\in\mathcal{O}(n)$

### v2

@@ A.3 Proof of Proposition 2
Proof of Theorem 4.1 Since $\mathcal{M}$ satisfies Assumption 1 and $h$ is smooth, we can take a neighborhood $\mathcal{B}_{p}\subseteq\mathbb{R}^{q}$ of $X$ such that $c_{1}$ and $h$ are $L_{c}$- and $L_{h}$-Lipschitz on $\mathcal{B}_{p}$, respectively, and $\operatorname{dist}({Y},\mathcal{M})\leq\rho_{\mathcal{M}}\|c_{1}({Y})\|_{2}$ holds for a constant $\rho_{\mathcal{M}}>0$ and any ${Y}\in\mathcal{B}_{p}$. Therefore, we can find a neighborhood $\mathcal{B}\subseteq\mathcal{B}_{p}$ such that given any $Y\in\mathcal{B}$, there exists $Y_{p}\in\mathcal{M}\cap\mathcal{B}_{p}$ such that $\|Y-Y_{p}\|_{2}=\operatorname{dist}(Y,\mathcal{M})$.
@@
where we employ the the triangle inequality $\|h(Y_{p})\|_{2}-\|h(Y)\|_{2}\leq\|h(Y_{p})-h(Y)\|_{2}$, and the Lipschitz continuity of $\phi$ and $h$. Therefore, setting $\rho=\rho_{\mathcal{M}}(1+L_{\phi}L_{h}C\rho_{\overline{\mathcal{K}}})+L_{\phi}C\rho_{\overline{\mathcal{K}}}$ concludes that $\operatorname{dist}(Y,\mathcal{M}\cap\mathcal{K})\leq\rho\|(c_{1}(Y),h(Y))\|_{2}$ for all $Y\in\mathcal{B}$, which together with the locally Lipschitz property of $(c_{1},h)$, verifies that $\mathcal{M}\cap\mathcal{K}$ satisfies Assumption 1. Applying Theorem 3.1 to $\mathcal{M}\cap\mathcal{K}=\{\tilde{X}\in\mathbb{R}^{q}\mid c_{1}(\tilde{X})=0,\,h(\tilde{X})=0\}$ gives
@@
$=\{\zeta\in\mathbb{R}^{q}\mid c_{1}^{\prime\prime}(X;\eta,\zeta)=0\}\cap\{\zeta\in\mathbb{R}^{q}\mid h^{\prime\prime}(X;\eta,\zeta)=0\}$
@@ C.3 $\mathcal{H}$ as the product of hyperbolic manifolds
$\overline{\mathcal{H}}=\phi_{\mathrm{LR}}^{-1}(\mathbb{H}^{n}_{m-1})=\{(L,R)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}\mid\bar{h}(L,R)=0\}\cap\mathcal{Z},$ (C.15)
@@ D.2.1 $\mathcal{U}$ as a Frobenius sphere
[…]
@@ D.3.1 $\mathcal{U}$ as an affine set
Appendix E Lemmas for section 5.2
This section is devoted to several technical lemmas essential for the analysis in section 5.2. We begin by relating the second-order tangent set to the second fundamental form on a smooth manifold in the following lemma.
Lemma 6
Given a smooth manifold $\overline{\mathcal{M}}$ embedded in an Euclidean space $\mathcal{W}$. Let $Y\in\overline{\mathcal{M}}$ and $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. The associated second-order tangent set can be identified as an affine subspace, $\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)=\mathrm{II}_{Y}(v,v)+\mathrm{T}_{\overline{\mathcal{M}}}(Y)$.
Proof
As shown in (lee2012manifolds, Proposition 5.16), the submanifold $\overline{\mathcal{M}}$ can be locally realized as the level set of a smooth submersion $h:\mathcal{W}\to\mathbb{R}^{k}$ (where $k$ is the codimension of $\overline{\mathcal{M}}$). That is, there exists a neighborhood $\mathcal{B}$ of $Y$ in $\mathcal{W}$ such that $\overline{\mathcal{M}}\cap\mathcal{B}=\{\tilde{Y}\in\mathcal{B}\mid h(\tilde{Y})=0\}$. Additionally, it holds that $\mathrm{T}_{\overline{\mathcal{M}}}(Y)=\operatorname{ker}(\mathrm{D}h_{Y})$.
It is revealed in (levin2025effect, formula (3.7)) that the second-order tangent set to the manifold can be characterized by $\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)=\{u_{v}\in\mathcal{W}\mid\mathrm{D}h_{Y}[u_{v}]+\mathrm{D}^{2}h_{Y}[v,v]=0\}$. That is, $\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)$ consists of all vectors $u_{v}\in\mathcal{W}$ satisfying the following linear equation:
$\mathrm{D}h_{Y}[u_{v}]=-\mathrm{D}^{2}h_{Y}[v,v],$ (E.1)
the solution set of which is an affine subspace given by the sum of a particular solution and the kernel of the linear operator $\mathrm{D}h_{Y}$. Note that $\operatorname{ker}(\mathrm{D}h_{Y})=\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, and thus it suffices to show that $\mathrm{II}_{Y}(v,v)$ is a particular solution of (E.1).
Let $\beta(t)$ be a smooth curve on $\overline{\mathcal{M}}$ such that $\beta(0)=Y$, $\beta^{\prime}(0)=v$, and we denote $w=\beta^{\prime\prime}(0)$. By definition of the second fundamental form, we have $\mathrm{II}(v,v)=\mathcal{P}_{\mathrm{N}_{\overline{\mathcal{M}}}(Y)}(w)$. Moreover, since the curve lies on the manifold, we have the identity $h(\beta(t))\equiv 0$. Twice differentiating it with respect to $t$ yields
$\mathrm{D}h_{\beta(t)}[\beta^{\prime\prime}(t)]+\mathrm{D}^{2}h_{\beta(t)}[\beta^{\prime}(t),\beta^{\prime}(t)]=0.$
Evaluating at $t=0$ and substituting $(\beta(0),\beta^{\prime}(0),\beta^{\prime\prime}(0))=(Y,v,w)$, we obtain $\mathrm{D}h_{Y}[w]=-\mathrm{D}^{2}h_{Y}[v,v]$. Substituting $w=\mathrm{II}_{Y}(v,v)+\mathcal{P}_{\mathrm{T}_{\overline{\mathcal{M}}}(Y)}(w)$ and recalling $\operatorname{ker}(\mathrm{D}h_{Y})=\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ lead to $\mathrm{D}h_{Y}[\mathrm{II}_{Y}(v,v)]=-\mathrm{D}^{2}h_{Y}[v,v]$, which concludes that $\mathrm{II}_{Y}(v,v)$ is a particular solution of (E.1).
The subsequent two lemmas examine the properties of the set $\mathcal{Q}_{Y}(v)$ defined in (5.5). Specifically, Lemma 7 characterizes the structure of $\mathcal{Q}_{Y}(0)$ for the case $v=0$, while Lemma 8 demonstrates the role of $\mathcal{Q}_{Y}(v)$ in optimization landscapes for general tangent vectors $v$.
Lemma 7
The set $\mathcal{Q}_{Y}(0)$ is a cone, i.e., if $\zeta\in\mathcal{Q}_{Y}(0)$, then $c\zeta\in\mathcal{Q}_{Y}(0)$ for all $c\geq 0$. Moreover, it holds that $\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathcal{Q}_{Y}(0)$.
Proof
Let $\zeta\in\mathcal{Q}_{Y}(0)$ and $c\geq 0$. By definition of $\mathcal{Q}_{Y}(0)$ in (5.5), there exist $\{v_{i}\}_{i\in\mathbb{N}}\subset\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ with $\mathrm{\bf L}_{Y}(v_{i})\to 0$, $u_{i}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v_{i})$, and $\{z_{i}\}_{i\in\mathbb{N}}\subseteq\operatorname{im}(\mathrm{\bf L}_{Y})$ such that $\zeta=\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}}\right)$. We consider the characterization of the second-order tangent set: $\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v)=\mathrm{II}_{Y}(v,v)+\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Thus, for each $i$, we can decompose $u_{i}$ as $u_{i}=\mathrm{II}_{Y}(v_{i},v_{i})+w_{i}$ for some $w_{i}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Then, we can introduce the scaled sequences $\tilde{v}_{i}:=\sqrt{c}v_{i}$ and $\tilde{u}_{i}:=cu_{i}$, which directly yields $\mathrm{\bf L}_{Y}(\tilde{v}_{i})=\sqrt{c}\mathrm{\bf L}_{Y}(v_{i})\to 0$. In addition, we have
$\tilde{u}_{i}=c\mathrm{II}_{Y}(v_{i},v_{i})+cw_{i}=\mathrm{II}_{Y}(\sqrt{c}v_{i},\sqrt{c}v_{i})+cw_{i}=\mathrm{II}_{Y}(\tilde{v}_{i},\tilde{v}_{i})+cw_{i},$
revealing that $\tilde{u}_{i}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;\tilde{v}_{i})$. Finally, setting $\tilde{z}_{i}=cz_{i}\in\operatorname{im}(\mathrm{\bf L}_{Y})$, the limit becomes $\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,\tilde{v}_{i}}(\tilde{u}_{i})+\tilde{z}_{i}}\right)=c\lim_{i\to\infty}\left({\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}}\right)=c\zeta\in\mathcal{Q}_{Y}(0)$.
Moreover, we can take $v_{i}=0$, which implies that $\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v_{i})=\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ (giorgi2010overviewsectengent, Proposition 9). Therefore, for any $u\in\operatorname{im}(\mathrm{\bf L}_{Y})$, set $u_{i}=0$, and $z_{i}=u$ to obtain $\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}=u$ for all $i\in\mathbb{N}$, indicating that $u=\lim_{i\to\infty}\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}\in\mathcal{Q}_{Y}(0)$. The arbitrariness of $u\in\operatorname{im}(\mathrm{\bf L}_{Y})$ concludes that $\operatorname{im}(\mathrm{\bf L}_{Y})\subseteq\mathcal{Q}_{Y}(0)$.
Lemma 8
Given a smooth function $f$ defined on $\mathcal{E}$ and a parameterization $(\overline{\mathcal{M}},\phi)$ for $\mathcal{M}$, suppose that $Y\in\overline{\mathcal{M}}$ is a second-order stationary point for problem (P-M) and denote $X=\phi(Y)$. Then for any $v\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, it holds that
$\langle{\nabla f(X),\zeta}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle\geq 0,\ \ \text{for all}\ \zeta\in\mathcal{Q}_{Y}(v).$
Proof
The first-order stationarity of $Y$ reveals that $\langle{\nabla\bar{f}(Y),v^{\prime}}\rangle=0$ for all $v^{\prime}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Substituting $\nabla\bar{f}(Y)=\mathrm{D}\phi_{Y}^{*}(\nabla f(X))$ shows that $\nabla f(X)\perp\operatorname{im}(\mathrm{\bf L}_{Y})$. According to the definition of $\mathcal{Q}_{Y}(v)$, any element $\zeta\in\mathcal{Q}_{Y}(v)$ can be realized by $\zeta=\lim_{i\to\infty}\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}$ with $\{v_{i}\}_{i\in\mathbb{N}}\subseteq\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, $\{z_{i}\}_{i\in\mathbb{N}}\subseteq\operatorname{im}(\mathrm{\bf L}_{Y})$, and the associated $u_{i}\in\mathrm{T}^{2}_{\overline{\mathcal{M}}}(Y;v_{i})$ satisfying $\mathrm{\bf L}_{Y}(v_{i})\to\mathrm{\bf L}_{Y}(v)$. Since $Y$ is second-order stationary for (P-M), the condition (5.1) indicates that for any $(v_{i},u_{i})$, it holds that $\langle{\nabla\bar{f}(Y),u_{i}}\rangle+\langle{v_{i},\nabla^{2}\bar{f}(Y)[v_{i}]}\rangle\geq 0$. Applying the equality (5.4), and substituting $z_{i}\in\operatorname{im}(\mathrm{\bf L}_{Y})$ together with the first-order condition $\nabla f(X)\bot\operatorname{im}(\mathrm{\bf L}_{Y})$ yield that $\langle{\nabla f(X),\mathrm{\bf Q}_{Y,v_{i}}(u_{i})+z_{i}}\rangle+\left\langle\mathrm{\bf L}_{Y}(v_{i}),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v_{i})]\right\rangle\geq 0$. Letting $i\to\infty$ in the inequality, we obtain $\langle{\nabla f(X),\zeta}\rangle+\left\langle\mathrm{\bf L}_{Y}(v),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v)]\right\rangle\geq 0$.
The next lemma investigates the property of $\varGamma_{Y}$ constructed in (5.5).
Lemma 9
Given a smooth function $f$ defined on $\mathcal{E}$ and a parameterization $(\overline{\mathcal{M}},\phi)$ for $\mathcal{M}$, suppose that $Y\in\overline{\mathcal{M}}$ is a second-order stationary point for problem (P-M), and $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$ with $X\!=\phi(Y)$. Then for any $\tau\in\varGamma_{Y}$, it holds that $\langle{\nabla f(X),\tau}\rangle=0$.
Proof
Recalling the definition of $\varGamma_{Y}$ (5.5), we denote $\tau\in\varGamma_{Y}$ by
$\tau=\mathcal{P}_{\hat{\mathrm{N}}_{\mathcal{M}}(X)}\left({\mathrm{D}\phi_{Y}(\mathrm{II}(v_{0},v_{1}))+\mathrm{D}^{2}\phi_{Y}[v_{0},v_{1}]}\right)\ \text{for some}\ v_{0}\in\ker(\mathrm{\bf L}_{Y}),\,v_{1}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y).$
Define a symmetric bilinear function on $\mathrm{T}_{\overline{\mathcal{M}}}(Y)\times\mathrm{T}_{\overline{\mathcal{M}}}(Y)$ as follows,
$q(\tilde{v}_{0},\tilde{v}_{1}):=\langle\nabla f(X),\mathrm{D}\phi_{Y}(\mathrm{II}(\tilde{v}_{0},\tilde{v}_{1}))+\mathrm{D}^{2}\phi_{Y}[\tilde{v}_{0},\tilde{v}_{1}]\rangle+\langle{\mathrm{\bf L}_{Y}(\tilde{v}_{0}),\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(\tilde{v}_{1})]}\rangle.$
Since $v_{0}\in\ker(\mathrm{\bf L}_{Y})$, we have $\mathrm{\bf L}_{Y}(v_{0})=0$, and thus
$\langle{\nabla f(X),\tau}\rangle=\langle{\nabla f(X),\mathcal{P}_{\hat{\mathrm{N}}_{\mathcal{M}}(X)}\left({\mathrm{D}\phi_{Y}(\mathrm{II}(v_{0},v_{1}))+\mathrm{D}^{2}\phi_{Y}[v_{0},v_{1}]}\right)}\rangle=q(v_{0},v_{1}).$
The last equality comes from the first-order optimality $\nabla\bar{f}(Y)=\mathrm{D}\phi_{Y}^{*}(\nabla f(X))\perp\mathrm{T}_{\overline{\mathcal{M}}}(Y)$, which, together with $\operatorname{im}(\mathrm{\bf L}_{Y})=\mathrm{T}_{\mathcal{M}}(X)$, reveals that $\nabla f(X)\in\hat{\mathrm{N}}_{\mathcal{M}}(X)$. Then, to prove $\langle{\nabla f(X),\tau}\rangle=0$, it suffices to prove $q(v_{0},v_{1})=0$ for all $v_{0}\in\ker(\mathrm{\bf L}_{Y}),\,v_{1}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. To see this, we note from (5.2) and (5.4) that $q(\tilde{v},\tilde{v})=\langle{\nabla\bar{f}(Y),u_{\tilde{v}}}\rangle+\langle{\tilde{v},\nabla^{2}\bar{f}(Y)[\tilde{v}]}\rangle$, and thus the second-order stationarity of $Y$ reveals that $q(\tilde{v},\tilde{v})\geq 0$ for all $\tilde{v}\in\mathrm{T}_{\overline{\mathcal{M}}}(Y)$. Considering the direction $v_{0}+tv_{1}$, we have $q(v_{0}+tv_{1},v_{0}+tv_{1})\geq 0$ for all $t$, which can be expanded as follows,
$q(v_{0}+tv_{1},v_{0}+tv_{1})=q(v_{0},v_{0})+2tq(v_{0},v_{1})+t^{2}q(v_{1},v_{1})\geq 0.$ (E.2)
Based on $\mathrm{\bf L}_{Y}(v_{0})=0$, it holds that $\mathrm{D}\phi_{Y}(\mathrm{II}(v_{0},v_{0}))+\mathrm{D}^{2}\phi_{Y}[v_{0},v_{0}]\in\operatorname{im}(\mathrm{\bf Q}_{Y,v_{0}})\subseteq\mathrm{T}^{2}_{\mathcal{M}}(X;0)=\mathrm{T}_{\mathcal{M}}(X)$, where the equality holds from (giorgi2010overviewsectengent, Proposition 9). Therefore, we have
$q(v_{0},v_{0})\!=\!\underbrace{\langle\nabla f(X),\mathrm{D}\phi_{Y}(\mathrm{II}(v_{0},v_{0})+\mathrm{D}^{2}\phi_{Y}[v_{0},v_{0}]\rangle}_{=0,\ \text{since}\ \nabla f(X)\perp\mathrm{T}_{\mathcal{M}}(X)}+\langle{\underbrace{\mathrm{\bf L}_{Y}(v_{0})}_{=0},\nabla^{2}f(X)[\mathrm{\bf L}_{Y}(v_{0})]}\rangle\!=0,$
which enforces $q(v_{0},v_{1})\geq 0$ to guarantee (E.2). Replacing $v_{1}$ with $v_{1}$ and repeating the analysis show that $q(v_{0},v_{1})=0$.
Appendix F Proof of Lemma 3
@@
$U_{i}\ \tilde{U}_{i\bot}\ \breve{U}_{i\bot}\ \tilde{U}_{iY\bot}\ U_{iY}]\in\mathcal{O}(m),\ \ \ \text{and}\ \ \ [V_{i}\ \tilde{V}_{i\bot}\ \breve{V}_{i\bot}\ \tilde{V}_{iY\bot}\ V_{iY}]\in\mathcal{O}(n)$